Documenta Mathematica

Extra Volume: Andrei A. Suslin’s Sixtieth Birthday, 2010

Preface  1

Denis Benois

Infinitesimal Deformations and the l-Invariant  5–31

Mikhail V. Bondarko Motivically Functorial Coniveau Spectral Sequences;

Direct Summands of Cohomology of Function Fields  33–117

Manuel Breuning and David Burns

On Equivariant Dedekind Zeta-Functions at s = 1  119–146

V. Chernousov Variations on a Theme of Groups Splitting by a Quadratic Extension and Grothendieck-Serre Conjecture for

Group Schemes F_{4} with Trivial g_{3} Invariant  147–169

Vincent Franjou and Wilberd van der Kallen

Power Reductivity over an Arbitrary Base  171–195

Eric M. Friedlander and Julia Pevtsova Generalized Support Varieties

for Finite Group Schemes  197–222

Thomas Geisser

On Suslin’s Singular Homology and Cohomology  223–249

Detlev W. Hoffmann Dimensions of Anisotropic

Indefinite Quadratic Forms II  251–265

Kevin Hutchinson, Liqun Tao Homology Stability for the Special Linear Group of a Field

and Milnor-Witt K-theory  267–315

Bruno Kahn Cohomological Approaches to SK_{1}

and SK_{2} of Central Simple Algebras  317–369

Nikita A. Karpenko with an Appendix by Jean-Pierre Tignol

Hyperbolicity of Orthogonal Involutions  371–392

Marc Levine

Slices and Transfers  393–443

Aurel Meyer, Zinovy Reichstein Some Consequences of

the Karpenko-Merkurjev Theorem  445–457

M. V. Nori and V. Srinivas

K-Theory and the Enriched Tits Building  459–513

Ivan Panin and Konstantin Pimenov Rationally Isotropic Quadratic Spaces

Are Locally Isotropic: II  515–523

S. Saito and K. Sato A p-adic Regulator Map and Finiteness Results

for Arithmetic Schemes  525–594

Michael Spieß

Twists of Drinfeld–Stuhler Modular Varieties  595–654

V. Valtman and S. Vostokov Artin-Hasse Functions

and their Invertions in Local Fields  655–660

Alexander Vishik

Rationality of Integral Cycles  661–670

Vladimir Voevodsky

Cancellation Theorem  671–685

Go Yamashita Bounds for the Dimensions

of p-Adic Multiple L-Value Spaces  687–723